A probability density function of a discrete random variable x is given by

f(x)={(1/2)x x=1,2,3,... 0 elsewhere}

use pgf to find a)E(x) b)V(x)




1
Expert's answer
2021-11-29T19:13:56-0500
"\displaystyle\sum_{n=1}^{\infin}q^n=\dfrac{q}{1-q}, |q|<1"

Differentiate both sides of the equality with respect to "q"


"\displaystyle\sum_{n=1}^{\infin}\dfrac{nq^n}{q}=\dfrac{1}{(1-q)^2}"

"\displaystyle\sum_{n=1}^{\infin}nq^n=\dfrac{q}{(1-q)^2}"

Similarly

Differentiate both sides of the equality with respect to "q"


"\displaystyle\sum_{n=1}^{\infin}\dfrac{n^2q^n}{q}=\dfrac{1}{(1-q)^2}+\dfrac{2q}{(1-q)^3}"


"\displaystyle\sum_{n=1}^{\infin}n^2q^n=\dfrac{q}{(1-q)^2}+\dfrac{2q^2}{(1-q)^3}"

a)


"E(X)=\displaystyle\sum_{x=1}^{\infin}x(\dfrac{1}{2})^x""q=1\/2, n=x"

"E(X)=\displaystyle\sum_{x=1}^{\infin}x(\dfrac{1}{2})^x=\dfrac{1/2}{(1-1/2)^2}=2"

b)


"E(X^2)=\displaystyle\sum_{x=1}^{\infin}x^2(\dfrac{1}{2})^x""q=1\/2, n=x"


"E(X^2)=\displaystyle\sum_{x=1}^{\infin}x^2(\dfrac{1}{2})^x=\dfrac{1/2}{(1-1/2)^2}+\dfrac{2(1/2)^2}{(1-1/2)^3}=6"


"Var(X)=E(X^2)-(E(X^2))^2=6-(2)^2=2"